What are the "zeros" of f(x) = 3x^3 - x^2 - 2x?
Thank you.

Answer by Cromlix(4381)

Hi there,
3x^3 - x^2 - 2x = 0
x(3x^2 - x - 2) = 0
x(3x + 2)(x - 1) = 0
x = 0
....
3x + 2 = 0
3x = -2
x = -2/3
......
x - 1 = 0
x = 1
....
Hope this helps :-)